OpenStudy (anonymous):

hi! where did i get wrong? 1-logbase4 x= -logbase8 x answer is supposed to be x=64? what i did... i made it 1=logbase4 x-logbase8 x then i changed the base thing... so i used log base 2 so i had... logbase2 x over 2 (from logbase2 4) - logbase2x over 3 (from logbase2 8) i factored out logbase2x so i had logbase2x (1/2-1/3) then i made it exponential so x/6=2 so i got 12????

OpenStudy (anonymous):

@hartnn please? :D

hartnn (hartnn):

how did u change the base ?

hartnn (hartnn):

\(\huge \log_b a = \frac{log a}{log b}\)

hartnn (hartnn):

oh wait you are correct till logbase2x (1/2-1/3)=1 so (1/6) log_2 x = 1 log_2 x =6 x= 2^6 = 64

hartnn (hartnn):

got the error ?

hartnn (hartnn):

@dydlf

OpenStudy (anonymous):

sorry for the late reply! anyway OHHHHHH i get it!! YES YES YES! thank youuu!!!

hartnn (hartnn):

welcome ^_^